OpenStudy (anonymous):

Rewrite the given equation in slope-intercept form: 15x - 3y = 12

OpenStudy (anonymous):

Determine the slope and the y-intercept for the linear equation: y = 6x

OpenStudy (ash2326):

Slope intercept form \[y=mx+c\] m=slope of the line c= y-intercept we have \[15x-3y=12\] Add 3y to both the sides, what would you get?

OpenStudy (anonymous):

i have no idea? :(

OpenStudy (ash2326):

\[15x-3y=12\] Add 3y to both the sides \[15x-3y+3y=12+3y\] \[15x-\cancel {3y}+\cancel{3y}=12+3y\] \[15x=12+3y\] Do you get this?

OpenStudy (anonymous):

y = 5x+4?

OpenStudy (ash2326):

close, \[15x-12=3y\] Divide both sides by 3

OpenStudy (anonymous):

So that is right. y = 5x+4

OpenStudy (anonymous):

oh 5x-4

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

let me try this one... Choose the slope and y-intercept of the equation: y = - 7x + 4

OpenStudy (ash2326):

yes :)

OpenStudy (anonymous):

um help...

OpenStudy (ash2326):

\[y=mx+c\] m= slope c= y intercept

OpenStudy (anonymous):

-7 (0,4)

OpenStudy (ash2326):

Correct :D

OpenStudy (anonymous):

Yea!! Determine the slope and the y-intercept for the linear equation: y = 6x

OpenStudy (ash2326):

I'll rewrite it for you \[y=6x+0\] Try now

OpenStudy (anonymous):

The slope is 6 and the y-intercept is (0, 0)

OpenStudy (ash2326):

yes ;D

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Determine the slope of a line perpendicular to a line whose slope is - 3.

OpenStudy (ash2326):

@smartrob If you have more questions, close this and ask a new. I'll answer this product of two perpendicular lines' slopes is -1 \[m1\times m2=-1\] here you have one line's slope

OpenStudy (anonymous):

ok